A particle of mass m is thrown vertically upward with velocity v0. The air resistance is mg cv^2 where c is a constant and v is the velocity at any time t. Show that the time taken by
the particle to reach the highest point is given by v0 underroot c= tan (gt underroot c )
1
Expert's answer
2020-03-03T15:06:32-0500

A particle thrown vertically with air resistance.

Initial speed of a particle: "v_0" (up). 

Air resistance modeled as "F_{air}=-cv^2," where "c" is a constant and "v" is the velocity at any time "t."

Using Newton’s 2nd law gives:


"m \text{\"{y}}=-mg-c\\text{\\.y}^2"

Alternatively, one can write


"{dv \over dt}=-g-{c \over m}v^2"

"dt=-{dv \over g+{c \over m}v^2}"

Integrate


"\displaystyle\int_{0}^{t_h}dt=\displaystyle\int_{v_0}^0-{dv \over g+{c \over m}v^2}"

"t_h=-\sqrt{{m \over cg}}\bigg[\arctan\big(\sqrt{{c \over mg}}v\big)\bigg]\begin{matrix}
 0 \\
 v_0
\end{matrix}"

"t_h=\sqrt{{m \over cg}}\arctan\big(\sqrt{{c \over mg}}v_0\big)"